# A Survey of Large Language Models in Medicine: Progress, Application, and Challenge

### v4

@@ B.2 Text Simplification
Text simplification is to generate a new text that recovers almost all the information of the original text while improving its readability. The task has potential in biomedical education, since one major characteristic of medical education is its opaque vocabulary [182, 289].
@@
Table 2 shows the commonly used text simplification datasets. Patel et al. [290] used NER to identify the medical terms. Lexical substitution is then applied to reduce the complexity of the text. Jeblick et al. [291] tested the performance of ChatGPT on simplifying self-collected radiology reports. They tried different prompting texts and found that the prompt ‘Explain this medical report to a child using simple language’ performs the best. However, by evaluating from a) factual correctness, b) completeness, and c) harmfulness, they concluded that GPT models may generate harmful texts which is unacceptable in the medical domain. Joseph et al. [76] evaluated the performance of zero-shot GPT-3 and Flan-T5 on MultiCochrane. They also fine-tuned mT5 [292] and Flan-T5 on this dataset. Yang et al. [293] introduced a data augmentation method for text simplification based on LLMs. For text without its simplified counterpart, they used GPT-3 to generate multiple choices for simplified text, and further trained a BERT model as the score to choose the best one as the simplification. Van et al. [81] transferred the simplification task into a prediction task. They assume they have the original text $(d_{1},d_{2},\cdots,d_{n})$ and a unfinished simplified text $(s_{1},s_{2},\cdots,s_{i})$. The task is to predict the next token $s_{i+1}$. In particular, at each time, the model receive $(d_{1},d_{2},\cdots,d_{n},s_{1},s_{2},\cdots,s_{i})$ as input and predict $s_{i+1}$. They used different models like BERT, RoBERTa, XLNet, and GPT-2. They also tried to combine the predicted token $s_{i+1}$ of all four models to outperform any of them.
@@
For evaluation, we use the bilingual evaluation understudy (BLEU) score [294]. We compare the performance of GPT-3 [76], Flan-T5 [76], and mT5 [76] on MultiCochrane (English) [76] (please see Table B.1). For AutoMeTS dataset [81], we use the accuracy of the next token $s_{i+1}$ as we discussed and we compare the performance of BERT, RoBERTa [81], XLNet [81], GPT-2 [81], and their combination [81] (please see Table B.1). The citations in this paragraph correspond to the sources that provide data on the performance of the models. For MultiCochrane, T5 models have much better performance than other baselines. For AutoMeTS, a combination of different models actually outperforms any of them, but overall, none of those models have a high accuracy or BLEU score.
@@ B.3 Text Generation
Text generation is a broad task that includes or relates to multiple specific tasks such as question answering (Sec A.1) and text summarization (Sec B.1). Here we focus on data-to-text tasks with open (instead of well-defined) answers, which involves taking structured data (e.g. a table) and producing text as output that describes these data. For example, generating patient clinic letters, radiology reports, and medical notes [295].
@@
Yermakov et al. [296] introduced a new dataset BioLeaflets and evaluated the performance of multiple LLMs on the data-to-text generation task. They found that T5 [297] and BARD [87] achieve the best performance for this task. However, multiple questions remain. Current LLMs may generate texts with typos, hallucinations, and repetitious words. Also, LLMs are not mature enough to produce coherent long text so far. Ranjit et al. [298] studied the task of generating reports for chest X-ray images. They also investigated and discussed the hallucinations that occurred in the GPT-generated texts. There are concerns about using model-generated pseudo-text for attacking since humans without expert knowledge cannot easily see the factual errors in the generated text. Rodriguez et al. [299] works on preventing attacks in the biomedical domain.

### v5

@@ B.2 Text Simplification
Text simplification is to generate a new text that recovers almost all the information of the original text while improving its readability. The task has potential in biomedical education, since one major characteristic of medical education is its opaque vocabulary [ondov2022survey, van2019evaluating].
@@
Table 4 shows the commonly used text simplification datasets. Patel et al. [patel2022summarization] used NER to identify the medical terms. Lexical substitution is then applied to reduce the complexity of the text. Jeblick et al. [jeblick2022chatgpt] tested the performance of ChatGPT on simplifying self-collected radiology reports. They tried different prompting texts and found that the prompt ‘Explain this medical report to a child using simple language’ performs the best. However, by evaluating from a) factual correctness, b) completeness, and c) harmfulness, they concluded that GPT models may generate harmful texts which is unacceptable in the medical domain. Joseph et al. [joseph2023multilingual] evaluated the performance of zero-shot GPT-3 and Flan-T5 on MultiCochrane. They also fine-tuned mT5 [xue2020mt5] and Flan-T5 on this dataset. Yang et al. [yang2023data] introduced a data augmentation method for text simplification based on LLMs. For text without its simplified counterpart, they used GPT-3 to generate multiple choices for simplified text, and further trained a BERT model as the score to choose the best one as the simplification. Van et al. [van2020automets] transferred the simplification task into a prediction task. They assume they have the original text $(d_{1},d_{2},\cdots,d_{n})$ and a unfinished simplified text $(s_{1},s_{2},\cdots,s_{i})$. The task is to predict the next token $s_{i+1}$. In particular, at each time, the model receive $(d_{1},d_{2},\cdots,d_{n},s_{1},s_{2},\cdots,s_{i})$ as input and predict $s_{i+1}$. They used different models like BERT, RoBERTa, XLNet, and GPT-2. They also tried to combine the predicted token $s_{i+1}$ of all four models to outperform any of them.
@@
For evaluation, we use the bilingual evaluation understudy (BLEU) score [papineni2002bleu]. We compare the performance of GPT-3 [joseph2023multilingual], Flan-T5 [joseph2023multilingual], and mT5 [joseph2023multilingual] on MultiCochrane (English) [joseph2023multilingual] (please see Table B.1). For AutoMeTS dataset [van2020automets], we use the accuracy of the next token $s_{i+1}$ as we discussed and we compare the performance of BERT, RoBERTa [van2020automets], XLNet [van2020automets], GPT-2 [van2020automets], and their combination [van2020automets] (please see Table B.1). The citations in this paragraph correspond to the sources that provide data on the performance of the models. For MultiCochrane, T5 models have much better performance than other baselines. For AutoMeTS, a combination of different models actually outperforms any of them, but overall, none of those models have a high accuracy or BLEU score.
@@ B.3 Text Generation
Text generation is a broad task that includes or relates to multiple specific tasks such as question answering (Sec A.1) and text summarization (Sec B.1). Here we focus on data-to-text tasks with open (instead of well-defined) answers, which involves taking structured data (e.g. a table) and producing text as output that describes these data. For example, generating patient clinic letters, radiology reports, and medical notes [liu2023utility].
@@
Yermakov et al. [yermakov2021biomedical] introduced a new dataset BioLeaflets and evaluated the performance of multiple LLMs on the data-to-text generation task. They found that T5 [T5] and BARD [manyika2023overview] achieve the best performance for this task. However, multiple questions remain. Current LLMs may generate texts with typos, hallucinations, and repetitious words. Also, LLMs are not mature enough to produce coherent long text so far. Ranjit et al. [ranjit2023retrieval] studied the task of generating reports for chest X-ray images. They also investigated and discussed the hallucinations that occurred in the GPT-generated texts. There are concerns about using model-generated pseudo-text for attacking since humans without expert knowledge cannot easily see the factual errors in the generated text. Rodriguez et al. [rodriguez2022cross] works on preventing attacks in the biomedical domain.